you have $50,000 to invest and two funds that you'd like to invest in the you risk it fund (fund y ) yield 14% interest the extra dull fund (fund x) yields 6% interest because of collage financial aid implication ,you don't think you can afford to earn more than $4,500 in interest income this year how much should you put in each fund ?
Answer by richwmiller(17219)

x+y=50000,
0.14*x+0.06*y=4500
x=50000-y
0.14*(50000-y)+0.06*y=4500
7000-0.14y+0.06*y=4500
-0.08*y=-2500
y=31250 at 6%
x=50000-y
x=18750 at 14%
check
0.14*18750+0.06*31250=4500
2625+1875=4500
4500=4500
ok